Construct a confidence interval for the population mean using a t-distribution: c = 99% m = 16.7 s = 3.2 n = 20.
Answer by Boreal(15235)

t df=19,0.995=2.86093
half-interval is t*s/sqrt(n)=2.861*3.2/sqrt(20)=2.047
this is added and subtracted to the mean for the CI.
(14.653, 18.747) or (14.6, 18.7)